BYD automotive must choose between two directions for the automaker’s future: developing fully electric-powered car or a hybrid-powered car. The company has $30 million capital available at a cost of capital of 10%. The $30 million capital needs to be paid back in 3.5 years.

A) The hybrid car project calls for investment of $10 million to build plant and assembly lines right now. For the next four years, the project cash flow from the project will be $5 million, $3 million, $3.5 million, and $3.5 million, respectively. After that, the project will be close down with no salvage value.

B) The electric car project calls for $20 million investment to purchase patents and another $10 million to build plant and assembly lines right now. For the next four years, the project cash flow from the project will be $15 million, $9 million, $9 million, and $11 million, respectively. After that, the project will be close down with no salvage value.

  1. Use the above information, fill the table of cash flow (careful about positive or negative sign of the cash flow) for each project.

Project Hybrid Car

Project Electric Car

Year

Cash Flow

Year

Cash Flow

0

0

1

1

2

2

3

3

4

4

  1. Based on the capital budgeting decision rules, calculate the Payback Period (PB), Discounted Payback Period (DPB), Net Present Value (NPV), and Profitability Index (PI) for each project.
  2. Use Excel function to calculate the Internal Rate of Return (IRR) for both projects.
Answer #1

B8 - fx =IRR(B3:37) B C D 1 Hybrid Car 2 year Cash flows pv@10% Present value Cumulative Cash flows Discounted Cumulative Cas

fx =IRR(13:17) N 18 H 1 Electric car 2 year Cash flows pv@10% $ (3,00,00,000.00) $ 1,50,00,000.00 2 $ 90,00,000.00 3 $ 90,00,